The piston diameter of a certain hand pump is 0.5 inch. The manager determines that the diameters are normally​ distributed, with a mean of 0.5 inch and a standard deviation of 0.004 inch. After recalibrating the production​ machine, the manager randomly selects 23 pistons and determines that the standard deviation is 0.0028 inch. Is there significant evidence for the manager to conclude that the standard deviation has decreased at the alpha equals 0.05 level of​ significance?

What are the correct hypotheses for this​ test?

Calculate the value of the test statistic.

Use technology to determine the​ P-value for the test statistic.

What is the correct conclusion at the α=0.05 level of​ significance?
Answer #1

Below are the null and alternative Hypothesis,
Null Hypothesis, H0: σ = 0.004
Alternative Hypothesis, Ha: σ < 0.004


Test statistic,
Χ^2 = (n-1)*s^2/σ^2
Χ^2 = (23 - 1)*0.0028^2/0.004^2
Χ^2 = 10.78

P-value Approach
P-value = 0.0223


As P-value < 0.05, reject the null hypothesis.